### 08 Apr 2026 v2

# Interventional Time Series Priors for Causal Foundation Models

Abstract
Prior-data fitted networks (PFNs) have emerged as powerful foundation models for tabular causal inference, yet their extension to time series remains limited by the absence of synthetic data generators that provide interventional targets. Existing time series benchmarks generate observational data with ground-truth causal graphs but lack the interventional data required for training causal foundation models. To address this, we propose CausalTimePrior, a principled framework for generating synthetic temporal structural causal models (TSCMs) with paired observational and interventional time series. Our prior supports configurable causal graph structures, nonlinear autoregressive mechanisms, regime-switching dynamics, and multiple intervention types (hard, soft, time-varying). We demonstrate that PFNs trained on CausalTimePrior can perform in-context causal effect estimation on held-out TSCMs, establishing a pathway toward foundation models for time series causal inference.
1 Introduction
Foundation models have transformed machine learning by enabling test-time inference without task-specific training. In tabular domains, prior-data fitted networks (PFNs) achieve this by pre-training transformers on synthetic datasets sampled from a prior distribution over data-generating processes (Müller et al., 2022; Hollmann et al., 2023). Recent work extends PFNs to causal inference: Do-PFN (Robertson et al., 2025) and CausalFM (Ma et al., 2025) train on synthetic structural causal models (SCMs) (Pearl, 2009) with interventional data, enabling in-context estimation of treatment effects from observational data alone.
However, extending causal PFNs to time series faces a fundamental obstacle: the lack of suitable synthetic data generators. While benchmarks like CausalTime (Cheng et al., 2024), TimeGraph (Ferdous et al., 2025), and CauseMe (Runge et al., 2019) provide time series with ground-truth causal graphs, they generate only observational data. Without interventional targets, one cannot train models to predict outcomes under interventions—the core task of causal inference.
We address this gap with CausalTimePrior [^1] https://github.com/thummd/CausalTimePrior, a framework for sampling temporal SCMs (TSCMs) together with paired observational and interventional time series (see Table 1). Our contributions are (1) a practical prior over discrete-time dynamic SCMs (Boeken and Mooij, 2024) that generates paired observational and interventional time series for training causal foundation models, (2) support for regime-switching TSCMs with Markov-driven structural breaks and interventional data—the first generator combining regime-switching dynamics (Balsells-Rodas et al., 2024) with interventional time series generation, and (3) preliminary experiments demonstrating that PFNs trained on CausalTimePrior can predict interventional outcomes given only observational context.
2 Related Work
Time series causal discovery benchmarks.
CausalTime (Cheng et al., 2024) fits neural networks to real observations and extracts causal graphs via importance analysis, producing realistic data with known ground truth. TimeGraph (Ferdous et al., 2025) generates synthetic time series from linear and nonlinear SCMs with configurable graph properties. CauseMe (Runge et al., 2019) benchmarks (including Lorenz-96 systems) and CausalRivers (Stein et al., 2025), the largest real-world benchmark (1,160 stations), offer physical ground-truth graphs. In contrast to CausalTimePrior, these existing methods are limited to observational data without interventions.
Generators with interventional support.
We identified only three frameworks supporting temporal interventions, each with limitations. CAnDOIT (Castri et al., 2024) generates time-lagged SCMs with hard interventions on known, single-node targets. It supports nonlinear mechanisms but only static intervention values. TECDI/RealTCD (Li et al., 2023; 2024) use structural VAR models with soft (TECDI) or hard (RealTCD) interventions. RealTCD handles unknown targets but is limited to linear mechanisms. CaTSG (Xia et al., 2025) implements do-calculus (Pearl et al., 2016) via diffusion models with a causal score function. While CaTSG implements interventions via learned diffusion models requiring training on specific datasets, CausalTimePrior generates interventional data analytically from explicit structural equations, enabling fast prior sampling without a separate generative model.
Table 1: Comparison of time series causal data generators. CausalTimePrior is the first to support regime-switching dynamics (changing causal structures over time).
Generator Interventions Nonlinear SCMs Time-varying Interventions Regime-switch
CausalTime
TimeGraph
CAnDOIT Hard
TECDI Soft
CaTSG do-calculus
CausalTimePrior All
Causal PFNs for tabular data.
Do-PFN (Robertson et al., 2025) pre-trains transformers on synthetic SCMs to predict conditional interventional distributions (CIDs) without knowing the causal graph. CausalFM (Ma et al., 2025) formalizes Bayesian priors over SCMs for back-door, front-door, and instrumental variable settings. Both demonstrate strong performance on i.i.d. tabular data but do not address temporal dependencies. Related work on counterfactual time series estimation includes CRN (Bica et al., 2020) and the Causal Transformer (Melnychuk et al., 2022), which estimate individualized treatment effects over time but require per-dataset training rather than in-context learning.
Time series foundation models.
Recent work has explored zero-shot forecasting via synthetic pre-training. ForecastPFN (Dooley et al., 2023) and TimePFN (Taga et al., 2025) pre-train transformers on synthetic data for multivariate forecasting. TempoPFN (Moroshan et al., 2025) pre-trains linear Recurrent Neural Networks (RNNs) for univariate forecasting, using diverse synthetic generators including Stochastic Differential Equations (SDEs), Gaussian processes, and causal kernels (CauKer) (Xie et al., 2025). While CauKer generator produces multivariate SCM-based time series, it lacks temporal lag structures and intervention support. These models target prediction rather than causal inference.
Regime-switching dynamics.
Switching Dynamical Systems (SDSs) with Markov Switching Models provide identifiability theory for regime-dependent causal discovery (Balsells-Rodas et al., 2024), but focus on inference from observational data without generating interventional datasets for training foundation models.
3 CausalTimePrior
We define a prior $\Pi$ over temporal structural causal models that generates paired observational and interventional time series suitable for training causal foundation models (see Algorithm 1).
3.1 Temporal Structural Causal Models
Following the Dynamic Structural Causal Model (DSCM) framework (Boeken and Mooij, 2024), we consider the discrete-time acyclic case. A temporal SCM (TSCM) $\mathcal{S}=(\mathcal{G},\mathbf{F},P_{\epsilon})$ consists of:
  • A time-lagged DAG $\mathcal{G}=(G_{0},G_{1},\ldots,G_{K})$ where $G_{0}\in\{0,1\}^{N\times N}$ encodes instantaneous (intra-slice) edges and $G_{k}$ encodes edges from time tkt-k to $t$ for lags $k\in\{1,\ldots,K\}$.
  • Structural equations $\mathbf{F}=\{f_{i}\}_{i=1}^{N}$ where:
    $X_{t}^{(i)}=f_{i}\left(\text{Pa}_{\mathcal{G}}(X_{t}^{(i)})\right)+\epsilon_{t}^{(i)},\quad\epsilon_{t}^{(i)}\sim P_{\epsilon}^{(i)}$ (1)
    with parents $\text{Pa}_{\mathcal{G}}(X_{t}^{(i)})=\{X_{t-k}^{(j)}:G_{k}[j,i]=1,k\in\{0,\ldots,K\}\}$.
3.2 Prior Distribution over TSCMs
Graph prior $\Pi_{\mathcal{G}}$.
We sample the number of variables $N\sim\text{Uniform}(3,N_{\max})$, maximum lag $K\sim\text{Uniform}(1,K_{\max})$, and edge probability $p\sim\text{Beta}(\alpha,\beta)$. Instantaneous edges $G_{0}$ are sampled from an Erdős-Rényi model (Erdős and Rényi, 1960) with acyclicity enforced via topological ordering. Lagged edges $G_{k}$ are sampled independently with probability decaying as $p\cdot\gamma^{k}$ for persistence factor $\gamma\in(0,1]$.
Mechanism prior $\Pi_{\mathbf{F}}$.
We sample mechanisms from multiple families. For simple mechanisms:
$f_{i}(\mathbf{x})=\sum_{j\in\text{Pa}(i)}w_{ij}\cdot\phi_{ij}(x_{j})+b_{i}$ (2)
where weights $w_{ij}\sim\mathcal{N}(0,\sigma_{w}^{2})$, biases $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2})$, and $\phi_{ij}$ is sampled uniformly from $\{\text{id},\sin,\cos,\tanh,|\cdot|,(\cdot)^{2},\exp(-|\cdot|)\}$. The diversity of activation functions ensures the prior covers a wide range of nonlinear temporal dynamics.
Noise prior $\Pi_{\epsilon}$.
Noise distributions are sampled per variable from $\{\mathcal{N}(0,\sigma^{2}),\text{Uniform}(-a,a),\text{Laplace}(0,b)\}$ with scale parameters drawn from suitable hyperpriors.
3.3 Intervention Types
Given a sampled TSCM $\mathcal{S}$, we generate interventional data by modifying structural equations (Eberhardt and Scheines, 2007). Let $I\subseteq\{1,\ldots,N\}$ denote intervention targets and $t_{I}\subseteq\{1,\ldots,T\}$ the intervention times. We provide examples for each intervention type in Appendix C.
Hard interventions.
(do-operator) Replace $X_{t}^{(i)}:=c$ for $i\in I,t\in t_{I}$, severing incoming edges.
Soft interventions.
Perturb the mechanism: $X_{t}^{(i)}=f_{i}(\text{Pa}(X_{t}^{(i)}))+\delta_{i}+\epsilon_{t}^{(i)}$ for shift $\delta_{i}\sim\mathcal{N}(\mu_{\delta},\sigma_{\delta}^{2})$.
Time-varying interventions.
Set $X_{t}^{(i)}:=c(t)$ where $c(t)$ follows a specified profile (step, ramp, sinusoidal, or sampled trajectory) (Hernán and Robins, 2020).
3.4 Data Generation Pipeline
For each training example, we:
  1. 1.
    Sample $\mathcal{S}\sim\Pi$ (graph, mechanisms, noise).
  2. 2.
    Sample intervention specification: targets $I$, times $t_{I}$, type, and values.
  3. 3.
    Generate observational series $\mathbf{X}^{\text{obs}}_{1:T}$ by forward simulation.
  4. 4.
    Generate interventional series $\mathbf{X}^{\text{int}}_{1:T}$ under $\text{do}(X^{(I)}_{t_{I}}=c)$.
  5. 5.
    Form training tuple: $(\mathbf{X}^{\text{obs}}_{1:T},I,t_{I},c,Y^{\text{int}}_{\tau})$ where $Y^{\text{int}}_{\tau}$ is the outcome variable at target time $\tau$ under intervention.
3.5 Regime-Switching Priors
Real-world time series often exhibit structural breaks where causal relationships change (Rahmani and others, 2025). We extend our prior to regime-switching TSCMs, following the Markov Switching Model framework (Balsells-Rodas et al., 2024):
$X_{t}^{(i)}=f_{i}^{(d_{t})}\left(\text{Pa}_{\mathcal{G}^{(d_{t})}}(X_{t}^{(i)})\right)+\epsilon_{t}^{(i)},\quad d_{t}\sim\text{Markov}(\mathbf{P})$ (3)
where $d_{t}\in\{1,\ldots,R\}$ indexes the active regime with transition matrix $\mathbf{P}$. Each regime has its own causal graph $\mathcal{G}^{(r)}$ and mechanisms $\mathbf{F}^{(r)}$. Regime transitions follow a sticky Markov chain ($P_{ii}\approx 0.9$) to model persistent causal structures. In our experiments, 15% of training TSCMs are regime-switching with $R\in\{2,3\}$ regimes. Combined with interventional data generation, this enables training PFNs that can reason about interventions under time-varying causal structures.
3.6 Design Choices and Assumptions
Complexity, identifiability, and causality.
[…]
Noise assumptions and prediction horizon.
Noise is additive and exogenous: $\epsilon_{t}^{(i)}$ is added after the nonlinear activation and sampled independently of parent values. It is also Markovian—i.i.d. across time—implying that the causal graph fully mediates all temporal dependencies; extending to non-Markovian (temporally correlated) noise would model latent confounders persisting across time and is an important future direction. At training time, the PFN queries outcomes 0–5 steps after intervention onset (70% downstream queries at 1–5 steps, 30% instantaneous), concentrating learning on the short-range causal propagation window; see Appendix B for further discussion.
4 Experiments
Prior Validation.
We validate CausalTimePrior by analyzing a dataset of 100K generated TSCMs. (1) Structural diversity: the prior generates TSCMs with $N\in[3,10]$ variables, $K\in[1,3]$ lags, and $T=50$ time steps, including 70% diverse nonlinear TSCMs, 15% chain structures, and 15% regime-switching TSCMs. Erdős-Rényi sampling with varied edge probabilities implicitly produces canonical causal motifs (confounders, mediators, colliders) as subgraphs. (2) Stability: 0% divergence rate (no NaN/Inf values) across all 100K samples, achieved through value clipping and careful mechanism parameterization. (3) Intervention coverage: hard, soft, and time-varying interventions with mean effect size of 17.98 (std 53.93), demonstrating substantial variability in intervention magnitudes across types. (4) Paired data quality: observational and interventional series maintain similar statistics (obs: $\mu=46.78$, $\sigma=242.85$; int: $\mu=41.56$, $\sigma=228.52$), confirming that interventions produce realistic counterfactual outcomes rather than out-of-distribution artifacts. Example visualization and full distributions of prior properties are shown in Appendix D and E.
Proof-of-Concept PFN.
As a preliminary demonstration, we train a simple 2-layer GRU-based PFN (128 hidden dim, 11 min on CPU) on 100K TSCMs from CausalTimePrior and evaluate on 1,000 held-out TSCMs. The model learns to distinguish causal from non-causal queries: Pred/GT ratio of 0.95 for intervened queries vs. 0.46 for non-causal queries (Table 2), and achieves comparable RMSE to per-dataset VAR baselines without per-sample fitting (Table 3). Implementation details, full results, baselines, a shuffled-intervention control experiment, ablations, generalizations, and an example are in Appendix F, G, H, I, and J.
5 Conclusion
CausalTimePrior addresses a critical gap in time series causal inference: the absence of synthetic generators with interventional data for training foundation models. By combining diverse temporal generators with principled intervention logic, it yields a prior over TSCMs with diverse intervention types. Our preliminary results suggest that PFNs trained on this prior can perform in-context causal effect estimation, opening a pathway toward foundation models for time series causality.
Limitations and future work.
The framework currently assumes Markovian noise and discrete-time dynamics; extensions to non-Markovian confounding and continuous-time processes are important future directions. Our Erdős-Rényi graph prior implicitly covers canonical causal structures (confounders, mediators, colliders) but does not explicitly stratify over them as Do-PFN does for tabular settings; adding structured temporal motifs could improve coverage. The prior has not been validated against real-world causal time series distributions. We plan to (1) scale training to larger models with explicit canonical structure sampling, (2) incorporate continuous-time dynamics, and (3) benchmark on semi-synthetic datasets derived from real observational data.
Acknowledgments
The authors are grateful for valuable discussions with Jake Roberston, Frank Hutter, and the Prior Labs team at the EurIPS’25 Workshop on AI for Tabular Data.
Appendix A CausalTimePrior Algorithm
Algorithm 1 formalizes the CausalTimePrior sampling procedure for generating paired observational and interventional time series from TSCMs.
Algorithm 1 CausalTimePrior Sampling
1:Input: Prior hyperparameters $\Pi=(\Pi_{\mathcal{G}},\Pi_{\mathbf{F}},\Pi_{\epsilon})$, sequence length $T$
2:Output: Observational series $\mathbf{X}^{\text{obs}}_{1:T}$, interventional series $\mathbf{X}^{\text{int}}_{1:T}$, intervention spec $I,t_{I},c$
3:
4: // Sample TSCM
[…]
7: Sample $G_{0}\in\{0,1\}^{N\times N}$ (acyclic via topological ordering)
8:for $k=1$ to $K$ do
9:  Sample $G_{k}$ with edge probability $p\cdot\gamma^{k}$
10:end for
[…]
12:
13:for $i=1$ to NN do
14:  Sample mechanism $f_{i}\sim\Pi_{\mathbf{F}}$ (nonlinear autoregressive)
15:  Sample noise $P_{\epsilon}^{(i)}\sim\Pi_{\epsilon}$
16:end for
[…]
18:
19: // Sample intervention specification
20: Sample targets $I\subseteq\{1,\ldots,N\}$
21: Sample times $t_{I}\subseteq\{1,\ldots,T\}$
22: Sample type $\in\{\text{hard},\text{soft},\text{time-varying}\}$
23: Sample value(s) $c$ or $c(t)$
24:
25: // Generate paired time series
[…]
28:
29:return $\mathbf{X}^{\text{obs}}_{1:T}$, $\mathbf{X}^{\text{int}}_{1:T}$, $(I,t_{I},c)$
Continuous-time extension.
CausalTimePrior currently generates discrete-time SCMs (Algorithm 1), but our autoregressive mechanisms have a natural continuous-time interpretation via the Euler-Maruyama discretization (Kloeden and Platen, 1992). Recent work on causal discovery in continuous-time SDEs (Manten et al., 2025) motivates extending CausalTimePrior to generate interventional data from SDE-based causal models. Consider a causal Ornstein-Uhlenbeck process $dX_{t}=\theta(\mu-X_{t})\,dt+\sigma_{w}\,dW_{t}$; applying Euler-Maruyama with step $\Delta t$ yields:
$x_{t+1}=\underbrace{(1-\theta\Delta t)}_{c_{2}}x_{t}+\underbrace{\theta\mu\Delta t}_{c_{1}}+\underbrace{\sigma_{w}\sqrt{\Delta t}}_{c_{3}}Z_{t},\quad Z_{t}\sim\mathcal{N}(0,1)$ (4)
which is precisely the AR(1) form our mechanism prior generates. This means each sampled discrete-time SCM can be viewed as an Euler-Maruyama discretization of a continuous-time causal SDE system (Thumm and Mijares, 2025). A natural extension is to sample continuous-time mechanisms directly—e.g., via Neural ODEs (Chen et al., 2018) or Neural SDEs (Kidger et al., 2021)—and discretize at variable time steps, enabling the prior to generate irregularly-sampled interventional time series.
Appendix B Prior Assumptions and Limitations
This section details the modeling assumptions underlying CausalTimePrior and their implications for identifiability and generalization.
Effect propagation and decay.
Causal effects in the prior decay through three mechanisms: (1) the maximum lag $K\in\{1,2,3\}$ bounds direct causal influence to at most 3 steps back, (2) geometric edge decay ($p_{k}=p\cdot 0.7^{k}$) makes distant lag connections increasingly sparse, and (3) bounded activations (e.g., $\tanh$, $\sin$) naturally dampen signal propagation. There is no explicit spectral radius constraint on the weight matrices; instead, stability is enforced empirically via value clipping to [1000,1000][-1000,1000] at each simulation step and divergence rejection (any trajectory with $|x|>10^{6}$ or NaN/Inf is discarded and resampled). A 50-step burn-in period is discarded to approximate the stationary distribution.
This design implies that causal effects attenuate rapidly—typically within 1–5 steps after intervention onset. While this is appropriate for systems with short-range temporal dependencies, it may underweight long-range causal effects (e.g., economic policy impacts over months). A PFN trained on this prior would inherit this short-range bias.
Markovian noise and identifiability.
The noise $\epsilon_{t}^{(i)}$ is sampled i.i.d. at each time step—independent across both time and variables. This Markovian assumption implies that the causal graph $\mathcal{G}$ fully mediates all temporal dependencies: there are no hidden common causes acting across time beyond what the lagged edges encode. Identifiability within the modeled lag window follows from the additive noise model (ANM) structure (Peters et al., 2014): each mechanism has the form $\sigma(\sum w_{j}x_{j}+b)+\epsilon$, and the use of non-Gaussian noise families (Uniform, Laplace) for non-root nodes supports distinguishing causal from non-causal associations under standard ANM identifiability results.
An extension to non-Markovian confounding—where $\text{Cov}(\epsilon_{t}^{(i)},\epsilon_{t-s}^{(j)})\neq 0$ for some s>0s>0—would model latent confounders that persist over time (e.g., unobserved trends or regime states not captured by the graph). This would break the assumption that interventional distributions depend only on the explicit causal structure and is an important direction for increasing the prior’s realism.
Interventional vs. counterfactual semantics.
CausalTimePrior generates interventional paired data, not counterfactual paired data. Concretely, when generate_pair produces an observational series $\mathbf{X}^{\text{obs}}$ and an interventional series $\mathbf{X}^{\text{int}}$, the two simulations draw independent noise realizations $\{\epsilon_{t}^{(i)}\}$ and $\{\tilde{\epsilon}_{t}^{(i)}\}$. This means the trajectories may differ even outside the intervention window—not because of any causal effect, but because they are driven by different exogenous noise.
This corresponds to the interventional query $P(\mathbf{X}\mid\text{do}(X^{(i)}_{t_{I}}=c))$: “what would a new draw from the system look like under this intervention?” The alternative is the counterfactual query $P(\mathbf{X}^{CF}\mid\mathbf{X}^{\text{obs}},\text{do}(X^{(i)}_{t_{I}}=c))$: “given this specific observational realization, what would have happened if we had intervened?” Counterfactual inference requires the three-step abduction-action-prediction procedure (Pearl, 2009): (1) infer the noise $\epsilon=\mathbf{X}^{\text{obs}}-f(\text{Pa})$ (abduction), (2) modify the structural equation for the intervened variable (action), and (3) re-simulate with the same noise (prediction).
In our framework, generating counterfactual pairs would require sharing the pre-sampled noise tensor between the observational and interventional simulation runs, so that the two trajectories are identical before the intervention onset and diverge only through causal propagation of the intervention effect. This is a natural extension that would enable training PFNs for counterfactual estimation—a strictly harder task than interventional prediction, since it requires reasoning about unit-level rather than population-level effects.
Noise model details.
All noise is additive (added after the nonlinear activation) and exogenous (independent of parent values). Root nodes (no parents in $G_{0}$) receive Gaussian noise with $\text{std}\sim\text{ShiftedExp}(\text{rate}=1.0,\text{shift}=0.1)$, providing larger driving noise. Non-root nodes receive noise from one of three variance-matched families (each with probability $\tfrac{1}{3}$): Gaussian $\mathcal{N}(0,\text{std}^{2})$, Uniform U(a,a)U(-a,a) with $a=\text{std}\sqrt{3}$, or Laplace $\text{Lap}(0,b)$ with $b=\text{std}/\sqrt{2}$, where $\text{std}\sim\text{ShiftedExp}(\text{rate}=10.0,\text{shift}=0.01)$. The smaller non-root noise ensures that most of a variable’s variance comes from its causal parents rather than its own noise term.
Stationarity scope.
Within a single SCM instance, mechanisms are stationary: weights, biases, activation functions, and noise distributions are fixed across all time steps. Non-stationarity enters only through regime-switching SCMs (15% of samples), where 2–3 regimes—each with its own causal graph and mechanisms—alternate according to a sticky Markov chain. The current framework does not support continuously time-varying coefficients, one-time structural breaks, or heteroscedastic noise. These extensions would broaden the prior’s coverage of real-world non-stationary dynamics.
PFN prediction horizon.
During training, 70% of queries target a different variable 1–5 steps after intervention onset (downstream causal effects), while 30% query the intervention target itself at the intervention time (instantaneous effects). The PFN architecture accepts arbitrary $(\text{intervention\_time},\text{query\_time})$ pairs as normalized inputs and can in principle predict at any horizon, but accuracy is expected to degrade beyond the 0–5 step training concentration. Extending the training distribution to longer horizons—potentially with curriculum learning (Bengio et al., 2009)—could improve long-range causal effect estimation.
Appendix C Intervention Type Examples
Figure 1 illustrates the three intervention types supported by CausalTimePrior, applied to the same 10-variable TSCM. Each row shows one type: (1) Hard intervention ($\text{do}(X_{0}:=3.0)$) replaces the variable’s mechanism entirely with a constant, severing all incoming causal edges during the intervention window—visible as a flat interventional trajectory. (2) Soft intervention ($\delta=1.5$) adds an additive shift to the mechanism output while preserving parental influence—the interventional trajectory tracks the observational one but is displaced. (3) Time-varying intervention ($\text{do}(X_{0}:=2\sin(2\pi t/L))$) replaces the mechanism with a time-dependent function, producing a sinusoidal interventional trajectory.
Figure 1: The three intervention types in CausalTimePrior. Left column: the intervention signal applied to $X_{0}$. Right column: resulting observational (blue) vs. interventional (red) trajectories. Yellow shading marks the intervention window $[30,70)$.
Figure 2 shows the four time-varying intervention profiles in detail, each applied to the same SCM and variable. The step profile jumps from 2-2 to +2+2 at the midpoint; the ramp linearly interpolates from 2-2 to +2+2; the sinusoidal profile follows $c(t)=2\sin(2\pi t/L)$; and the sampled trajectory draws independent $c(t)\sim\mathcal{N}(0,4)$ at each time step. The dashed line shows the intervention signal $c(t)$ on the secondary axis.
Figure 2: Four time-varying intervention profiles. Blue: observational trajectory. Red: interventional trajectory. Dashed: intervention signal $c(t)$ (right axis). All four profiles are applied to the same variable in the same TSCM.
Appendix D Example Visualizations
Figure 3 shows an example of paired observational and interventional time series generated from CausalTimePrior. The hard intervention on Variable 4 between $t=20$ and $t=80$ causes a clear divergence between the observational (blue) and interventional (red) trajectories during and after the intervention period. Figure 4 displays all 6 variables in the sampled TSCM, with Variable 4 (the intervention target) highlighted. The propagation of intervention effects through the causal graph is visible in downstream variables, while non-causally connected variables remain unaffected.
Figure 3: Paired observational and interventional time series for the intervention target variable. The yellow shaded region indicates the intervention period. The divergence between blue (observational) and red (interventional) trajectories demonstrates the causal effect of the intervention.
Figure 4: All variables in a sampled 6-variable TSCM with a hard intervention on Variable 4. The intervention target is highlighted with a yellow background. Causal effects propagate through the graph structure, affecting downstream variables while leaving non-causally connected variables unchanged.
Appendix E Prior Property Distributions
Figure 5 shows the distributions of key properties across 100K TSCMs sampled from CausalTimePrior. (a) Graph sizes are approximately uniform over $N\in[3,10]$. (b) Intervention types are split 50% hard, 30% soft, and 20% time-varying. (c) Intervention effect magnitudes span several orders of magnitude (median 1.4) on a log scale, ensuring the prior covers both subtle and large causal effects. (d) Intervention start times are concentrated in the second half of the sequence to allow sufficient observational context. (e) Edge probabilities follow a Beta(2,5) prior (mean 0.29), producing mostly sparse graphs. (f) Intervention values are approximately Gaussian-distributed around zero.
Figure 5: Distributions of prior properties across 100K sampled TSCMs from CausalTimePrior. The prior produces diverse graph structures, intervention types, and effect magnitudes.
Appendix F Implementation Details
We implement a simple proof-of-concept architecture using a 2-layer GRU encoder (128 hidden dim). The temporal encoder processes the observational time series $\mathbf{X}^{\text{obs}}_{1:T}$, the intervention encoder embeds the intervention specification $\text{do}(X^{(i)}_{t}=c)$ (which variable, when, what value), and the query encoder embeds the prediction query (which variable to predict, when). The prediction head combines these encodings to output $P(Y_{\tau}^{\text{int}}|\text{do}(X^{(i)}_{t}=c),\mathbf{X}^{\text{obs}}_{1:T})$ as a Gaussian distribution with predicted mean and standard deviation. The training objective is:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathcal{S}\sim\Pi}\mathbb{E}_{\mathbf{X}^{\text{obs}}_{1:T}\sim P_{\text{obs}}^{\mathcal{S}}}\mathbb{E}_{(x_{t},y_{\tau})\sim P_{\text{int}}^{\mathcal{S}}}\left[-\log q_{\theta}(y_{\tau}|\text{do}(x_{t}),\mathbf{X}^{\text{obs}}_{1:T})\right]$ (5)
where $q_{\theta}$ is a Gaussian with predicted mean and variance, $\mathbf{X}^{\text{obs}}_{1:T}$ is the observational time series context, $\text{do}(x_{t})$ is the temporal intervention query at time $t$, and $y_{\tau}$ is the interventional outcome at target time $\tau$.
Prior hyperparameters.
$N_{\max}=10$, $K_{\max}=3$, $\alpha=2$, $\beta=5$ (sparse graphs), $\gamma=0.7$ (lag decay), $\sigma_{w}=1.0$, $\sigma_{b}=0.5$.
Training.
Adam optimizer, learning rate $10^{-4}$, batch size 32, sequence length 50, 15 epochs on 100K TSCMs. Training takes approximately 11 minutes on CPU (Intel/AMD with AVX2). The model checkpoint requires  330KB of storage.
Architecture.
We use a simplified 2-layer GRU encoder with 128 hidden dimensions and Gaussian prediction head (mean + standard deviation). We chose a GRU over transformer architectures (as used in Do-PFN and TimePFN) for computational simplicity in this proof-of-concept: the GRU processes variable-length sequences efficiently and its recurrent state naturally captures temporal dependencies. This choice prioritizes fast iteration ($\sim$11 min training on CPU) over architectural optimality; scaling to transformer or GatedDeltaProduct architectures (Moroshan et al., 2025) is an important next step.
Implementation.
CausalTimePrior is implemented from scratch, drawing conceptual inspiration from TempoPFN’s diverse generator design and intervention logic from Do-PFN. The core is a TemporalSCM class that supports both sample_observational(T) and sample_interventional(T, intervention) methods, enabling paired data generation from the same underlying causal structure with time-lagged dependencies and multiple intervention types.
Appendix G Results
We train a 2-layer GRU encoder (128 hidden dim) for 15 epochs on 100K TSCMs and evaluate on 1,000 held-out TSCMs using three query types: (1) Intervened (query the intervention target itself), (2) Downstream (query a variable causally reachable from the intervention), and (3) Non-causal (query a variable with no causal path from the intervention).
[…]
Baselines.
We compare against a Vector Autoregression (VAR-OLS) fitted per-dataset, PCMCI+ (Runge, 2020) which discovers causal graphs via conditional independence tests, and a mean prediction baseline (Table 3). SimpleCausalPFN achieves comparable RMSE to VAR-OLS (176.4 vs 176.5) while requiring no per-dataset fitting. PCMCI+ achieves lower overall RMSE (161.4) by leveraging discovered causal structure, but requires expensive per-sample graph discovery.
Shuffled-intervention control.
To test whether the model’s query-type sensitivity reflects learned causal structure rather than distributional artifacts, we evaluate with randomly shuffled intervention targets. Under shuffling, predictions for intervened queries change substantially (+33% mean prediction) and Pred/GT degrades from 0.95 to 1.26, indicating the model is sensitive to intervention target information. However, non-causal predictions remain low (+13%), suggesting the model has also partially learned distributional properties of variable positions, motivating richer architectures that more explicitly encode causal graph structure.
Table 2: Three-way evaluation on held-out TSCMs. NMSE (Normalized MSE = MSE/Var(GT); NMSE<<1 is better than predicting the mean). The model’s Pred/GT ratio is highest for intervened queries (0.95) and lowest for non-causal queries (0.46), suggesting learned causal understanding. Overall NMSE $\approx$ 1.0 indicates limited absolute prediction quality.
Query Type Queries RMSE $\downarrow$ NMSE $\downarrow$ Mean Pred Mean GT Pred/GT $\uparrow\downarrow=1$
Intervened 573 226.39 1.41 33.91 35.68 0.95
Downstream 270 231.31 0.67 57.50 67.70 0.85
Non-causal 157 143.15 0.66 15.66 33.99 0.46
Overall 1000 216.87 0.99 37.41 44.06 0.85
Table 3: Comparison with baselines on held-out TSCMs. SimpleCausalPFN achieves comparable RMSE to VAR-OLS while requiring no per-dataset fitting. PCMCI+ achieves the lowest RMSE but requires per-sample causal graph discovery.
Method RMSE $\downarrow$ MAE $\downarrow$ Effect Dir. Acc. $\uparrow$ Effect Size Corr. $\uparrow$
Oracle 0.00 0.00 100.0% 1.000
PCMCI+ 161.38 28.67 74.1% 0.784
SimpleCausalPFN 176.45 34.92 70.4% 0.821
VAR-OLS 176.45 33.86 93.7% 0.821
Mean Prediction 256.18 73.26 60.4% 0.286
Appendix H Intervention Type Ablation
We investigate whether training on diverse intervention types (hard, soft, time-varying) improves performance compared to training only on hard interventions. Table 4 compares the mixed-intervention model (trained on 100K TSCMs with all intervention types) against a hard-only model (trained on 10K TSCMs with only hard interventions) on the same three-way test set.
Table 4: Intervention type ablation. The mixed-intervention model achieves higher effect direction accuracy and effect size correlation compared to the hard-only model.
Model RMSE $\downarrow$ MAE $\downarrow$ Effect Dir. Acc. $\uparrow$ Effect Size Corr. $\uparrow$
Mixed (100K) 216.87 81.68 70.4% 0.821
Hard-only (10K) 212.16 69.68 63.9% 0.691
The mixed model achieves higher effect direction accuracy (70.4% vs 63.9%) and effect size correlation (0.821 vs 0.691), suggesting that diversity in intervention types during training improves the model’s ability to reason about causal effects. While the hard-only model achieves slightly lower overall RMSE (212.16 vs 216.87), this is likely due to the simpler prediction task when only hard interventions are present.
Appendix I Out-of-Distribution Generalization
We evaluate whether the model generalizes to TSCMs with structural properties outside the training distribution. We generate an out-of-distribution (OOD) test set of 1,000 TSCMs with: (1) larger graphs ($N\in[8,10]$ vs training mean $\sim$6), (2) maximum lag $K=3$, (3) denser graphs (edge probability $\geq 0.3$ vs training mean $\sim$0.29), and (4) only complex nonlinear mechanisms (sin, cos, square, tanhReLU).
Table 5: Out-of-distribution generalization. Performance degrades on OOD TSCMs with larger, denser graphs and complex mechanisms, but the model retains basic causal structure (downstream RMSE >> intervened RMSE).
Test Set RMSE $\downarrow$ NMSE $\downarrow$ MAE $\downarrow$ Effect Dir. Acc. $\uparrow$ Effect Size Corr. $\uparrow$
In-distribution 216.87 0.99 81.68 70.4% 0.821
OOD 265.97 0.72 163.25 62.7% 0.599
As expected, performance degrades on OOD data: RMSE increases from 216.87 to 265.97 and effect size correlation drops from 0.821 to 0.599. The lower OOD NMSE (0.72 vs 0.99) reflects higher ground-truth variance in the OOD test set rather than better prediction quality. However, the model retains some causal understanding, with intervened queries (RMSE=237.01) outperforming downstream queries (RMSE=313.92), consistent with the in-distribution pattern.
Appendix J Causal vs. Correlation-Based Prediction
To illustrate how the PFN leverages causal structure rather than correlations, consider a concrete test case. In sample 626 from our test set, a 3-variable temporal TSCM is intervened on variable 2 at $t=25$, and we query variable 0 at $t=28$. There is no causal path from variable 2 to variable 0, but the observational time series exhibit a correlation of 0.49-0.49 between them. The ground truth interventional value (0.056-0.056) is nearly identical to the observational baseline (0.094-0.094), confirming a near-zero causal effect (0.039). The PFN predicts 0.050-0.050 (error $0.005$), correctly recognizing the absence of causal influence despite the spurious correlation. In contrast, VAR-OLS—which relies on learned correlations without distinguishing causal from non-causal associations—predicts 0.992-0.992 (error $0.936$), a $177\times$ larger prediction error. This pattern generalizes: across 157 non-causal queries, the PFN achieves lower prediction error than VAR-OLS in 45% of cases, with the largest gains precisely on samples exhibiting high spurious correlations ($|\rho|>0.3$).